You have ​$5800 to invest today at 11​% interest compounded annually.

a. Find how much you will have accumulated in the account at the end of​ (1) 6 ​years, (2) 12 ​years, and​ (3) 18 years.

b. Use your findings in part a to calculate the amount of interest earned in​ (1) the first 6 years​ (years 1 to 6​), ​(2) the second 6 years​ (years 7 to 12​), and​ (3) the third 6 years​ (years 13 to 18​). c. Compare and contrast your findings in part b.  

Explain why the amount of interest earned increases in each succeeding 6 dash year period.
Answer #1

a.

1.

Future Value at the end of Year 6 = 5,800(1.11)6 = $10,848.40

2.

Future Value at the end of Year 12 = 5,800(1.11)12 = $20,291.01

3.

Future Value at the end of Year 18 = 5,800(1.11)18 = $37,952.61

b.

Interest earned from Year 1 to Year 6 = 10,848.40 - 5,800 = $5,048.40

Interest earned from Year 6 to Year 12 = 20,291.01 - 10,848.40 = $9,442.61

Interest earned from Year 13 to Year 18 = 37,952.61 - 20,291.01 = $17,661.60